Soybean is 18%protein and cornmeal is 9% protein, how many lbs you mix for 360lbs at 16% protein?
Answer by mananth(16946)

Soyabean is 18%protein and cornmeal is 9% protein, how many lbs you mix for 360lbs at 16% protein?
soyabean---------------- cornmeal-------------- mix
18---------------------- 9---------------------16
let quantity of soyabean be x
quantity of cornmeal will be 360-x
0.18x+ 0.9*(360-x)= 360*0.16
0.18x+324x- 0.9x = 57.6
-0.78x= -266.4
x=341.54
cornmeal 18.46